OpenStudy (anonymous):

Describe the transformation: 2^x -->4^x

OpenStudy (accessdenied):

So, we have the starting function: f(x) = 2^x Somehow, we get to a function f(x) = 4^x. What I notice is that 2 and 4 are related in terms of 4 being the same as 2^2. Being an exponent, it seems like this could be useful to our function. Try rewriting 4 as 2^2 in the function and using some exponent properties to get a form of y = 2^(something). :)